please help me with this equation....x^2+6x+8=0
Found 2 solutions by richwmiller, kekegeter :
Answer by richwmiller(17219)

find factors of 8 that add up to 6
2 and 4 come to mind
(x+2)*(x+4)=0
x=-2 x=-4

Answer by kekegeter (97)

x^2+2x+4x+8=0
x[x^2/x+2x/x]+2^2[2^2*x/2^2+2^3/2^2]
x(x+2)+2^2(x+2)=0
(x+2)(x+4)=0
x+2=0
x+4=0
x=-2
x=-4